OpenStudy (anonymous):

w+1/2=1&6/8. what would w equal?

OpenStudy (australopithecus):

1 % 6/8 is the same as 1 + 6/8 Remember to add/subtract fractions they must have the same denominator, thus convert 1 to 8/8 you can do this because 8/8 = 1 w + 1/2 = 1 + 6/8 w + 1/2 = 8/8 + 6/8 w + 1/2 = (8 + 6)/8 w + 1/2 = 14/8 w = 14/8 - 1/2 now I'm going to reduce 14/8 by dividing the top and bottom by 2, remember as long as you do the same thing to the top and bottom of a fraction it doesn't change its value for example 8*2/8*2 = 8/8 = 8/3 / 8/3 = 8/8 = 1 w = 14/2 / 8/2 - 1/2 w = 7/4 - 1/2 now we need to change the denominator on one of these fractions in order to subtract them w = 7/4 - (1*2)/(2*2) w = 7/4 - 2/4 w = (7 - 2)/4 w = 5/4

OpenStudy (australopithecus):

I hope you follow

OpenStudy (anonymous):

um no, we r so confused, he is only in the 5th grade so we dont understand how u got that....... sorry for wasting your time. We dont understand thoo......

OpenStudy (australopithecus):

Not wasting my time, my time is worthless. What exactly don't you understand?

OpenStudy (anonymous):

instead of writing your question turn it into 1 and 6/8 - 1 and a half which would eaqual 1 and 1/4

OpenStudy (australopithecus):

show me the parts of my solution that dont make sense to you and I will try to explain them better

OpenStudy (anonymous):

Wait i just figured it out, thank you for your help, :D

OpenStudy (australopithecus):

a lot of this stuff you just have to convince yourself of. Play around with numbers to help clarify my explanation

OpenStudy (anonymous):

hey did i help???

OpenStudy (australopithecus):

sure why not

OpenStudy (anonymous):

yes u did, both of u did, :P